# Tensor networks for black hole interiors: non-isometries, quantum extremal surfaces, and wormholes

### v1

@@ 1 Introduction
This work takes one step towards a more realistic qudit model of the black hole interior. The model used in [1] for the semi-classical interior implemented a limited, binary notion of locality: you’re either behind the black hole horizon or you’re not. We aim to include a slightly more refined notion of geometry by allowing for relative positions of degrees of freedom in the interior and exterior of the black hole. Holographic tensor networks are a natural candidate to provide this notion of locality – using qudits, they describe bulk locality in the framework of error correcting codes [8, 9, 10]. Furthermore, the geometry provided by the network makes available tools such as Ryu-Takayanagi (RT)-like formulas for studying QESs and operator reconstruction within entanglement wedges.
Early tensor network descriptions of AdS/CFT pioneered by the HaPPY code [9] focused on regions exterior to any horizon. Built out of perfect tensors, these codes were designed to act as exact isometries in order to capture the quantum error correcting properties of AdS/CFT [8]. Black holes were represented in these early networks by excising tensors behind the horizon, leaving dangling legs representing fundamental black hole degrees of freedom along the horizon. Viewing these legs as inputs, the remaining annular-shaped tensor network would isometrically encode the fundamental black hole degrees of freedom on the horizon in the CFT degrees of freedom on the boundary. In this work, we will refer to such tensor network codes outside the horizon as the “outer code.”
However, tensor networks need not only be useful for the exterior of the black hole. An observer falling past the horizon should still be able to give a local description of the surroundings, even when they’ve passed into the interior. Thus the locality provided by a tensor network should also be at least approximately applicable to the interior of a black hole. The possibility of extending a tensor network into the interior is corroborated by the work of [10], which built holographic tensor networks from random tensors. By increasing the entanglement of the bulk state passed into the random tensor network, a phase transition in minimal surfaces consistent with the appearance of a black hole was observed. No tensors behind a horizon needed to be excised to observe this transition. This further suggests that tensor networks are useful for describing both the interior and the exterior.
@@ 4.1 Wormholes in dynamically generated states
Had we instead chosen to make the reservoir a second AdS space and allowed the radiation $R_{\text{out}}$ to collapse into a second AdS black hole, these new connections would eventually lead to a two-sided black hole. We note that TNs for two-sided black holes have been considered in the context of MERA networks [33] and the python’s lunch conjecture [20]. In both of these cases, the wormhole has more structure and is “longer” than our example, built out of its own tensors rather than just additional connections. This feature cannot be captured in our model, as we only describe evolution of states on top of an already existing tessellation. Dynamics of the tessellation itself could allow for the wormhole to grow in length. We leave exploration of these more sophisticated dynamics to future work.
@@ 6 Conclusion
This work takes another step forward in creating a more realistic qudit model of black hole interiors and understanding the importance of non-isometries in black hole codes. Additionally, we find that this model provides a useful way to understand features such as operator reconstruction and models of python’s lunch. We are hopeful that further improvements in the non-isometric codes initiated by PHEVA will continue to push our understanding of black hole information.
@@ Acknowledgements
We are grateful to Chris Akers for helpful discussions. The images of tessellations used in figures throughout this paper were generated using the KaleidoTile program. The authors are supported by the Department of Energy under grants DE-SC0010005 and DE-SC0020360.

### v2

@@ 1 Introduction
This work takes one step towards a more realistic qudit model of the black hole interior. The model used in [1] for the semi-classical interior implemented a limited, binary notion of locality: you’re either behind the black hole horizon or you’re not. We aim to include a slightly more refined notion of geometry by allowing for relative positions of degrees of freedom in the interior and exterior of the black hole. Holographic tensor networks are a natural candidate – using qudits, they describe bulk locality built from the quantum informational properties of the boundary. The geometry provided by such tensor network constructions makes available tools such as Ryu-Takayanagi (RT)-like formulas for studying QESs and operator reconstruction within entanglement wedges.
Tensor network descriptions of AdS/CFT were pioneered by [8], where a discrete bulk was built from $1d$ critical Ising models using MERA tensor networks. Later, the HaPPy code [9] was created to incorporate the error correcting properties of AdS/CFT [10] into a holographic tensor network. Built out of perfect tensors, the HaPPy code was designed to act as an exact isometry on bulk input states, encoding them redundantly in the boundary. These constructions focused on regions exterior to any horizon and represented black holes by excising tensors in their interior, leaving dangling legs representing fundamental black hole degrees of freedom along the horizon. Viewing these legs as inputs, the remaining annular-shaped tensor network would isometrically encode the fundamental black hole degrees of freedom on the horizon in the CFT degrees of freedom on the boundary. In this work, we will refer to such tensor network codes outside the horizon as the “outer code.”
However, tensor networks need not only be useful for the exterior of the black hole. An observer falling past the horizon should still be able to give a local description of the surroundings, even when they’ve passed into the interior. Thus the locality provided by a tensor network should also be at least approximately applicable to the interior of a black hole. A great deal of work has focused on the application of tensor networks to black hole interiors. For example, tensor networks have been used to study the interior of two-sided black holes [11], the complexity of black hole interiors [12], and in constructions of the python’s lunch proposal [13]. Furthermore, random tensor networks constructed in [14] were used to observe phase transitions in minimal surfaces consistent with the formation of a black hole. All of these examples suggest that the excised portions of a holographic tensor network behind a horizon can be filled in to give a local description of black hole interiors.
@@ 4.1 Wormholes in dynamically generated states
Had we instead chosen to make the reservoir a second AdS space and allowed the radiation $R_{\text{out}}$ to collapse into a second AdS black hole, these new connections would eventually lead to a two-sided black hole. We note that tensor networks for two-sided black holes have been considered in [11, 12, 13]; in all of these cases, the wormhole has more structure and is “longer” than our example, built out of its own tensors rather than just additional connections. This feature cannot be captured in our model, as we only describe evolution of states on top of an already existing tessellation. Dynamics of the tessellation itself could allow for the wormhole to grow in length. We leave exploration of these more sophisticated dynamics to future work.
@@ 6 Conclusion
As a final comment, we note that other works have considered other aspects of non-isometry and locality in holography. The work of [40] uses non-isometric maps to construct “overlapping qubits” that can be used to model a local effective field theory with a system that contains many fewer degrees of freedom. Here we have focused on locality for PHEVA-like holographic maps, and it would be interesting to further compare these results. We are hopeful that further improvements in these non-isometric codes for holography and black holes will continue to push our understanding of quantum gravity and black hole information.
@@ Acknowledgements
We are grateful to Chris Akers for helpful discussions, and to Brian Swingle and Wissam Chemissany for discussions on related work. The images of tessellations used in figures throughout this paper were generated using the KaleidoTile program. The authors are supported by the Department of Energy under grants DE-SC0010005 and DE-SC0020360.